# Weak-Lensing Shear-Selected Galaxy Clusters from the Hyper Suprime-Cam Subaru Strategic Program: I. Cluster Catalog, Selection Function and Mass–Observable Relation

### v1

@@ 3 Modeling Framework and Injection Simulations
A widely adopted analytical model for weak-lensing peak counts was proposed in Fan et al. (2010) and has been applied to constrain cosmology in Shan et al. (2012, 2018) and Liu et al. (2023). However, such a model heavily relies on the assumption that various sources of scatter on the weak-lensing observable are Gaussian in nature and cannot probe the full complexity of these systematic uncertainties. Lin & Kilbinger (2015) proposed a semi-analytical model by injecting halo into N-body simulation to quantify the full impact of projection from LSS, intrinsic alignment, and survey masks. In this work, we adopt a more sophisticated semi-analytical model to quantify the mass–observable relation. Instead of painting halos into N-body simulations, we inject the lensing signals of halos into the observed weak-lensing mass maps to determine $P(\nu|M,z,\boldsymbol{\mathbf{p}})$. This is achieved by injecting the synthetic shear distortions into the ellipticity of the source galaxies observed in the HSC-Y3 shape catalog.
@@
Here, $\hat{M_{\kappa}}$ and $\theta_{\mathrm{s}}$ are parameters representing the analytic aperture mass and the characteristic angular scale of the cluster, which will be precisely defined later in Sec. 3.3. These parameters capture the shape of the lensing profile. This way, we divide the mass–observable relation $P(\nu|M,z,\boldsymbol{\mathbf{p}})$ into two pieces: (1) $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ is the re-parametrized scaling relation that connects the shape of a lensing profile to the final observable $\nu$; and (2) $P(\hat{M_{\kappa}},\theta_{\mathrm{s}}|M,z,\boldsymbol{\mathbf{p}})$, which converts the cluster’s physical properties $\left(M,z\right)$ to the shape of the lensing profile parameterized by $\left(\hat{M_{\kappa}},\theta_{\mathrm{s}}\right)$. The first term, $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$, accounts for the measurement uncertainties including shape noise, cosmic shear, shape measurement bias, the variation in survey depth, etc. This needs to be determined through computationally expensive injection simulations. However, given a lensing profile determined by $\left(\hat{M_{\kappa}},\theta_{\mathrm{s}}\right)$, the process of injecting a synthetic cluster signal, re-detecting it on the resulting mass maps, and quantifying the relation $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ is insensitive to cosmological parameters and can be computed in advance. The cosmological dependency goes entirely into the conversion $P(\hat{M_{\kappa}},\theta_{\mathrm{s}}|M,z,\boldsymbol{\mathbf{p}})$. This term also captures the intrinsic uncertainties, such as the intrinsic scatter in the weak-lensing observable due to the diversity of halo profiles (e.g., halo concentration and triaxiality) and the weak-lensing mass bias arising from the inaccurate assumption about the halo profiles. The modeling of $P(\hat{M_{\kappa}},\theta_{\mathrm{s}}|M,z,\boldsymbol{\mathbf{p}})$ will be examined in-depth in our cosmological analysis (Chiu et al., 2024).
@@ 3.1 The Mock Cluster Samples
In practice, we sample 2554880 clusters for each connected field in our survey. These clusters are sampled uniformly in the log-parameter space with the boundaries of $5\times 10^{12}\,M_{\odot}h^{-1}\leq M_{200\mathrm{c}}\leq 5\times 10^{16}\,M_{\odot}h^{-1}$, $0.02\leq z_{\mathrm{cl}}\leq 2.00$, and $10^{-2}\,\mathrm{arcmin}\leq\theta_{\mathrm{s}}\leq 3\times 10\,\mathrm{arcmin}$.
@@
Figure 4: Examples of injecting cluster lensing signal into the HSC-Y1 XMM field shape catalog following the procedure outlined in Sec. 3.2. The left and the right panel each shows two cases that are accepted or rejected by step (v) in Sec. 3.2 in order to remove cases where we selected peaks that are not mainly associated with the injected halos. The dashed circles in these plots indicate the $5$ arc minute range around $\boldsymbol{\mathbf{x}}_{\mathbf{diff}}$ above which we reject a detection.
@@ 3.2 Injection Simulations
Fig. 4 shows four examples in our injection simulation following the procedure outlined above. For each example, we show the weak-lensing map before and after we inject the mock cluster lensing signal as described in step (iii) together with the difference map introduced in step (v). The right panel demonstrates two cases that are rejected in step (v) to avoid detecting peaks that are not mainly associated with the injected halos. The top row shows a rejected case as it fails to meet the $|\boldsymbol{\mathbf{x}}_{\mathrm{inj}}-\boldsymbol{\mathbf{x}}_{\mathrm{diff}}|\leq 5^{\prime}$ criterion. The bottom row shows another case that is rejected as it has $|\nu_{\mathrm{inj}}-\nu_{\mathrm{diff}}|>\Delta_{s}$ if we choose $\Delta_{s}=3.0$. This does not imply that the detected peak here is completely unassociated with the injected halo. The same case would be considered as a detection if we allow for a larger ranging of scattering and choose $\Delta_{s}=4.0$ for instance. Therefore, the allowed range of scattering $\Delta_{s}$ cannot be determined through injection simulation alone and must be calibrated by leveraging external information. In Sec. 4.3, we discuss how $\Delta_{s}$ affects the number count of weak lensing peaks predicted by our framework and leverage optical counterparts to inform $\Delta_{s}$. In our cosmological analysis (Chiu et al., 2024), the scatter parameter $\Delta_{s}$ will be self-calibrated by the number count $N(\nu)$ of the shear-selected clusters.
@@
Figure 5: Examples of injecting a single mock halo at various locations on the weak lensing mass map derived from the HSC-Y1 data. The singal-to-noise ratio $\nu_{\mathrm{inj}}$ after each injection is given in the middle panel. Here, not detected indicates a peak is not found on the map after we inject the halo, most likely due to masking; and rejected indicates a peak is found but does not satisfy the requirement given in step (v) in Sec. 3.2 for us to determine that it is associated with the injected halo. The rightmost two panels show the distributions of peak signal-to-noise ratio $\nu$ and the distance between the observed peak and the halo center from 5000 injections of this same halo. As a comparison, the distribution of peak signal-to-noise ratio predicted from an uniform shape noise is indicated as the dashed blue curve.
@@ 4.1 Results from Injection Simulations
Fig 5 shows the results of injecting the same cluster into different locations on the weak-lensing mass map. An example of the weak-lensing mass maps before and after the injection is presented in the leftmost panel. Meanwhile, the central panel shows that the same cluster is detected with various signal-to-noise ratios $\nu$ at different locations, revealing significant scatter due to the variations in both the signal from large-scale structures and the shape noise observed across the survey footprint. By performing the injections repeatedly on different locations on the maps, we are able to fully capture the measurement uncertainties in $\nu$ directly from the data without any assumption about the nature of these uncertainties. In some cases, the cluster is not detected due to the survey masking; while in other cases, a peak is found but is rejected due to the criterion in eq. 19. This means that either the peak has large angular offset to the injected center ($\arcmin>5\arcmin$), or the signal-to-noise ratio $\nu$ of the peak is significantly scattered beyond the range we consider ($|\nu_{\mathrm{inj}}-\nu_{\mathrm{diff}}|>\Delta_{s}$).
In the rightmost panels in Fig 5, we show the distribution of the signal-to-noise ratio (top panel) and the offset between the detected peak and the injection location (bottom panel). In the top panel, we also show the distribution of $\nu$ assuming only the presence of an uniform and Gaussian shape noise (blue dashed line) as a comparison. This is obtained by a normal distribution centered at $\hat{M_{\kappa}}(M,z,c)/\sigma$ with $\sigma$ evaluated as (van Waerbeke, 2000; Mandelbaum et al., 2018b)
@@ 4.2 Completeness Function and Scaling Relation
Fig. 6 shows the selection function and scaling relation derived from injection simulations performed over the HSC-Y1 XMM field with the selection $\nu_{\mathrm{min}}=4.7$. The leftmost panel shows the selection function $\mathcal{C}(\hat{M_{\kappa}},\theta_{\mathrm{s}})$; the middle panel shows the scaling relation $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ at three different points on the $\hat{M_{\kappa}}$$\theta_{\mathrm{s}}$ plane (indicated as crosses in the leftmost panel); and the rightmost panel shows the scatter of $\nu$ as a function of $\hat{M_{\kappa}}$, with each dot indicating the result from one injection. In the leftmost panel, we observe that the completeness function does not reach unity even for the most massive halos. As discussed before, this is because we inject halos into an area that is larger than the actual survey footprint. The completeness function saturates at the effective coverage of the survey footprint with respect to the injection area.
Similar to Fig. 5, the dashed histograms in the middle panel show the distributions of $\nu$ at different $\left(\hat{M_{\kappa}},\theta_{\mathrm{s}}\right)$ assuming only the effect of an uniform shape noise. We observe that the scatter in $\nu$ from the injection-based results is larger than those of the shape-noise-only distributions, regardless of the quantities $\left(\hat{M_{\kappa}},\theta_{\mathrm{s}}\right)$. This is more clearly seen in the rightmost panel, where we show the 68.45% confidence interval of the distribution $P(\nu|\hat{M_{\kappa}})$ as a function of $\hat{M_{\kappa}}$ derived from injection simulations (blue) and from analytical calculations assuming only the shape noise (green). At the high-$\hat{M_{\kappa}}$ end, halos that have been partly obscured by the survey boundaries or masks could still produce significant peaks to be detected, allowing more down-scatter; at the low-$\hat{M_{\kappa}}$ end, significantly stronger up-scatter is predicted by the injection simulations due to the presence of large-scale structures.
@@ 4.3 Validations
Following eq. (9) and (10), the number counts of shear-selected clusters as a function of the signal-to-noise ratio $\nu$ are calculated as
@@
The remaining component yet to be specified is $P(\hat{M_{\kappa}},\theta_{\mathrm{s}}|M,z,\boldsymbol{\mathbf{p}})$, the cosmology-dependent conversion from halos’ physical properties $(M,z)$ to their lensing properties $\left(M_{\kappa},\theta_{\mathrm{s}}\right)$. This term will be comprehensively modeled in our cosmological analysis (Chiu et al., 2024) to self-consistently take into account the intrinsic uncertainties associated with the halo. For the purpose of validation, we adopt a simplified modeling framework and write
@@
where $P(\hat{M_{\kappa}}|\theta_{\mathrm{s}},M_{\mathrm{WL}},z)$ is a Dirac delta function centered at $\hat{M_{\kappa}}(M_{\mathrm{WL}},z,\theta_{\mathrm{s}})$ given by Eq. (24), and $P(\theta_{\mathrm{s}}|M_{\mathrm{WL}},z)$ links the angular size of a halo to its weak lensing mass $M_{\mathrm{WL}}$at a given redshift, which we model through the mass–concentration relation. Here, we assume that the mass–concentration relation follows a log-normal distribution around the mean value predicted by Diemer & Joyce (2019) with a scatter $\sigma_{\ln c}=0.30$.
@@
Figure 7: Validation test on the invariance of the scaling relation $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ derived under different reference cosmologies. The three black lines (solid, dashed, dash-dotted) show the number counts predicted for the HSC-Y1 XMM field using $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ derived under three different sets of cosmological parameters while the dotted blue line is the prediction considering only an uniform shape noise. These results are compared to the number counts in a mock cluster samples (orange circles) and the number counts of real observed peak (green squares) in this field. The mock sample is constructed through drawing a Poisson realization of the halo mass function and directly inject them into the HSC-Y1 shape catalog to determine which of them can be observed (see Sec. 4.3.2). For visualization purpose, the green data points are shifted to the right by 0.05 to avoid overlap.
@@ 4.3.1 Cosmological Independence
Here, $\boldsymbol{\mathbf{p}}$ is fixed to the Planck cosmology to compute terms that are known to depend on cosmology, while $\boldsymbol{\mathbf{p}}_{i}$ represents the three sets of cosmological parameters given in Table 1 that is used to compute $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$. The number counts predicted by the three different $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}},\boldsymbol{\mathbf{p}}_{i})$ are shown in Fig. 7. As can be seen, the resulting cluster number counts $\mathrm{d}N_{i}/\mathrm{d}\nu$ are in excellent agreement with one another, strongly demonstrating that the scaling relation $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ is insensitive to the underlying cosmology. This offers great computational advantages as it allows us to compute $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ in advance and we will only need to compute the other cosmology-dependent terms in eq. (LABEL:eq:number_count_model) in the cosmological analysis.

### v2

@@ 3 Modeling Framework and Injection Simulations
A widely adopted analytical model for weak-lensing peak counts was proposed in Fan et al. (2010) and has been applied to constrain cosmology in Shan et al. (2012, 2014, 2018) and Liu et al. (2023). However, such a model heavily relies on the assumption that various sources of scatter on the weak-lensing observable are Gaussian in nature and cannot probe the full complexity of these systematic uncertainties. Lin & Kilbinger (2015) proposed a semi-analytical model by injecting halo into N-body simulation to quantify the full impact of projection from LSS, intrinsic alignment, and survey masks. In this work, we adopt a more sophisticated semi-analytical model to quantify the mass–observable relation. Instead of painting halos into N-body simulations, we inject the lensing signals of halos into the observed weak-lensing mass maps to determine $P(\nu|M,z,\boldsymbol{\mathbf{p}})$. This is achieved by injecting the synthetic shear distortions into the ellipticity of the source galaxies observed in the HSC-Y3 shape catalog.
@@
Here, $\hat{M_{\kappa}}$ and $\theta_{\mathrm{s}}$ are parameters representing the analytic aperture mass and the characteristic angular scale of the cluster, which will be precisely defined later in Sec. 3.3. These parameters capture the shape and amplitude of the lensing profile. This way, we divide the mass–observable relation $P(\nu|M,z,\boldsymbol{\mathbf{p}})$ into two pieces: (1) $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ is the re-parametrized scaling relation that connects the shape and amplitude of a lensing profile to the final observable $\nu$; and (2) $P(\hat{M_{\kappa}},\theta_{\mathrm{s}}|M,z,\boldsymbol{\mathbf{p}})$, which converts the cluster’s physical properties $\left(M,z\right)$ to the shape and amplitude of the lensing profile parameterized by $\left(\hat{M_{\kappa}},\theta_{\mathrm{s}}\right)$. The first term, $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$, accounts for the measurement uncertainties including shape noise, cosmic shear, shape measurement bias, the variation in survey depth, etc. This needs to be determined through computationally expensive injection simulations. However, given a lensing profile determined by $\left(\hat{M_{\kappa}},\theta_{\mathrm{s}}\right)$, the process of injecting a synthetic cluster signal, re-detecting it on the resulting mass maps, and quantifying the relation $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ is insensitive to cosmological parameters and can be computed in advance. The cosmological dependency goes entirely into the conversion $P(\hat{M_{\kappa}},\theta_{\mathrm{s}}|M,z,\boldsymbol{\mathbf{p}})$. This term also captures the intrinsic uncertainties, such as the intrinsic scatter in the weak-lensing observable due to the diversity of halo profiles (e.g., halo concentration and triaxiality) and the weak-lensing mass bias arising from the inaccurate assumption about the halo profiles. The modeling of $P(\hat{M_{\kappa}},\theta_{\mathrm{s}}|M,z,\boldsymbol{\mathbf{p}})$ will be examined in-depth in our cosmological analysis (Chiu et al., 2024).
@@ 3.1 The Mock Cluster Samples
In practice, we sample 2554880 clusters for each connected field in our survey. These clusters are sampled uniformly in the $\log\left[M_{200\mathrm{c}}\right]$$z_{\mathrm{cl}}$$\log\left[\theta_{\mathrm{s}}\right]$ parameter space with the boundaries of $5\times 10^{12}\,M_{\odot}h^{-1}\leq M_{200\mathrm{c}}\leq 5\times 10^{16}\,M_{\odot}h^{-1}$, $0.02\leq z_{\mathrm{cl}}\leq 2.00$, and $0.01^{\prime}\leq\theta_{\mathrm{s}}\leq 30^{\prime}$.
@@
Figure 4: Examples of injecting cluster lensing signal into the HSC-Y1 XMM field shape catalog following the procedure outlined in Sec. 3.2. The left and the right panel each shows two cases that are accepted or rejected by step (v) in Sec. 3.2 in order to remove cases where we select peaks that are not mainly associated with the injected halos. The dashed circles in these plots indicate the $5$ arc minute range around $\boldsymbol{\mathbf{x}}_{\mathbf{diff}}$ beyond which we reject a detection.
@@ 3.2 Injection Simulations
Figure 4 shows four examples in our injection simulation following the procedure outlined above. For each example, we show the weak-lensing map before and after we inject the mock cluster lensing signal as described in step (iii) together with the difference map introduced in step (v). The right panel demonstrates two cases that are rejected in step (v) to avoid detecting peaks that are not mainly associated with the injected halos. The top row shows a rejected case as it fails to meet the $|\boldsymbol{\mathbf{x}}_{\mathrm{inj}}-\boldsymbol{\mathbf{x}}_{\mathrm{diff}}|\leq 5^{\prime}$ criterion. The bottom row shows another case that is rejected as it has $|\nu_{\mathrm{inj}}-\nu_{\mathrm{diff}}|>\Delta_{s}$ if we choose $\Delta_{s}=3.0$. This does not imply that the detected peak here is completely unassociated with the injected halo. The same case would be considered as a detection if we allow for a larger ranging of scattering and choose $\Delta_{s}=4.0$ for instance. Therefore, the allowed range of scattering $\Delta_{s}$ cannot be determined through injection simulation alone and must be calibrated by leveraging external information. In Sec. 4.3, we discuss how $\Delta_{s}$ affects the number count of weak lensing peaks predicted by our framework and leverage optical counterparts to inform $\Delta_{s}$. In our cosmological analysis (Chiu et al., 2024), the scatter parameter $\Delta_{s}$ will be self-calibrated by the number count $N(\nu)$ of the shear-selected clusters.
@@
Figure 5: Examples of injecting a single mock halo at various locations on the weak lensing mass map derived from the HSC-Y1 data. The peak singal-to-noise ratio $\nu_{\mathrm{inj}}$ after each injection is given in the middle panel. Here, not detected indicates a peak is not found on the map after we inject the halo, most likely due to masking; and rejected indicates a peak is found but does not satisfy the requirement given in step (v) in Sec. 3.2 for us to determine that it is associated with the injected halo. The rightmost two panels show the distributions of peak signal-to-noise ratio $\nu$ and the distance between the observed peak and the halo center from 5000 injections of this same halo. As a comparison, the distribution of peak signal-to-noise ratio predicted from an uniform shape noise is indicated as the dashed blue curve.
@@ 4.1 Results from Injection Simulations
Figure 5 shows the results of injecting the same cluster into different locations on the weak-lensing mass map. An example of the weak-lensing mass maps before and after the injection is presented in the leftmost panel. Meanwhile, the central panel shows that the same cluster is detected with various signal-to-noise ratios $\nu$ at different locations, revealing significant scatter due to the variations in both the signal from large-scale structures and the shape noise observed across the survey footprint. By performing the injections repeatedly on different locations on the maps, we are able to fully capture the measurement uncertainties in $\nu$ directly from the data without any assumption about the nature of these uncertainties. In some cases, the cluster is not detected due to the survey masking; while in other cases, a peak is found but is rejected due to the criterion in Equation (19). This means that either the peak has large angular offset to the injected center ($\arcmin>5\arcmin$), or the signal-to-noise ratio $\nu$ of the peak is significantly scattered beyond the range we consider ($|\nu_{\mathrm{inj}}-\nu_{\mathrm{diff}}|>\Delta_{s}$).
In the rightmost panels in Figure 5, we show the distribution of the signal-to-noise ratio (top panel) and the offset between the detected peak and the injection location (bottom panel). In the top panel, we also show the distribution of $\nu$ assuming only the presence of an uniform and Gaussian shape noise (blue dashed line) as a comparison. This is obtained by a normal distribution centered at $\hat{M_{\kappa}}(M,z,c)/\sigma$ with $\sigma$ evaluated as (van Waerbeke, 2000; Mandelbaum et al., 2018b)
@@ 4.2 Completeness Function and Scaling Relation
Figure 6 shows the selection function and scaling relation derived from injection simulations performed over the HSC-Y1 XMM field with the selection $\nu_{\mathrm{min}}=4.7$. The leftmost panel shows the selection function $\mathcal{C}(\hat{M_{\kappa}},\theta_{\mathrm{s}})$; the middle panel shows the scaling relation $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ at three different points on the $\hat{M_{\kappa}}$$\theta_{\mathrm{s}}$ plane (indicated as crosses in the leftmost panel); and the rightmost panel shows the scatter of $\nu$ as a function of $\hat{M_{\kappa}}$, with each dot indicating the result from one injection. In the leftmost panel, we observe that the completeness function does not reach unity even for the most massive halos. As discussed before, this is because we inject halos into an area that is larger than the actual survey footprint. The completeness function saturates at the effective coverage of the survey footprint with respect to the injection area.
Similar to Figure 5, the dashed histograms in the middle panel show the distributions of $\nu$ at different $\left(\hat{M_{\kappa}},\theta_{\mathrm{s}}\right)$ assuming only the effect of an uniform shape noise. We observe that the scatter in $\nu$ from the injection-based results is larger than those of the shape-noise-only distributions, regardless of the quantities $\left(\hat{M_{\kappa}},\theta_{\mathrm{s}}\right)$. This is more clearly seen in the rightmost panel, where we show the 68.45% confidence interval of the distribution $P(\nu|\hat{M_{\kappa}})$ as a function of $\hat{M_{\kappa}}$ derived from injection simulations (blue) and from analytical calculations assuming only the shape noise (green). At the high-$\hat{M_{\kappa}}$ end, halos that have been partly obscured by the survey boundaries or masks could still produce significant peaks to be detected, allowing more down-scatter; at the low-$\hat{M_{\kappa}}$ end, significantly stronger up-scatter is predicted by the injection simulations due to the presence of large-scale structures.
@@ 4.3 Validations
Following Equation (9) and (10), the number counts of shear-selected clusters as a function of the signal-to-noise ratio $\nu$ are calculated as
@@
The last component to be specified is $P(\hat{M_{\kappa}},\theta_{\mathrm{s}}|M,z,\boldsymbol{\mathbf{p}})$, the cosmology-dependent conversion from halos’ physical properties $(M,z)$ to their lensing properties $\left(M_{\kappa},\theta_{\mathrm{s}}\right)$. This term will be comprehensively modeled in our cosmological analysis (Chiu et al., 2024) to self-consistently take into account the intrinsic uncertainties associated with the halo. For the purpose of validation, we adopt a simplified modeling framework and write
@@
where $P(\hat{M_{\kappa}}|\theta_{\mathrm{s}},M_{\mathrm{WL}},z)$ is a Dirac delta function centered at $\hat{M_{\kappa}}(M_{\mathrm{WL}},z,\theta_{\mathrm{s}})$ given by Equation (24), and $P(\theta_{\mathrm{s}}|M_{\mathrm{WL}},z)$ links the angular size of a halo to its weak lensing mass $M_{\mathrm{WL}}$at a given redshift, which we model through the mass–concentration relation. Here, we assume that the mass–concentration relation follows a log-normal distribution around the mean value predicted by Diemer & Joyce (2019) with a scatter $\sigma_{\ln c}=0.30$.
@@
Figure 7: Validation test on the invariance of the scaling relation $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ derived under different reference cosmologies. The three black lines (solid, dashed, dash-dotted) show the number counts predicted for the HSC-Y1 XMM field using $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ derived under three different sets of cosmological parameters while the dotted blue line is the prediction considering only an uniform shape noise. These results are compared to the number counts in a mock cluster samples (orange circles) and the number counts of real observed peaks (green squares) in this field. The mock sample is constructed through drawing a Poisson realization of the halo mass function and directly inject them into the HSC-Y1 shape catalog to determine which of them can be observed (see Sec. 4.3.2). For visualization purpose, the green data points are shifted to the right by 0.05 to avoid overlap.
@@ 4.3.1 Cosmological Independence
Here, $\boldsymbol{\mathbf{p}}$ is fixed to the Planck cosmology to compute terms that are known to depend on cosmology, while $\boldsymbol{\mathbf{p}}_{i}$ represents the three sets of cosmological parameters given in Table 1 that is used to compute $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$. The number counts predicted by the three different $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}},\boldsymbol{\mathbf{p}}_{i})$ are shown in Figure 7. As can be seen, the resulting cluster number counts $\mathrm{d}N_{i}/\mathrm{d}\nu$ are in excellent agreement with one another, strongly demonstrating that the scaling relation $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ is insensitive to the underlying cosmology. This offers great computational advantages as it allows us to compute $P(\nu|\hat{M_{\kappa}},\theta_{\mathrm{s}})$ in advance and we will only need to compute the other cosmology-dependent terms in Equation (LABEL:eq:number_count_model) in the cosmological analysis.